# Reinforcing Image Generation with Collaborative Semantic-level and Token-level CoT

### v1

@@ Abstract
Recent advancements in large language models have demonstrated how chain-of-thought (CoT) and reinforcement learning (RL) can improve performance. However, applying such reasoning strategies to the visual generation domain remains largely unexplored. In this paper, we present T2I-R1, a novel reasoning-enhanced text-to-image generation model, powered by RL with a bi-level CoT reasoning process. Specifically, we identify two levels of CoT that can be utilized to enhance different stages of generation: (1) the semantic-level CoT for high-level planning of the prompt and (2) the token-level CoT for low-level pixel processing during patch-by-patch generation. To better coordinate these two levels of CoT, we introduce BiCoT-GRPO with an ensemble of generation rewards, which seamlessly optimizes both generation CoTs within the same training step. By applying our reasoning strategies to the baseline model, Janus-Pro, we achieve superior performance with 13% improvement on T2I-CompBench and 19% improvement on the WISE benchmark, even surpassing the state-of-the-art model FLUX.1. Code is available at: https://github.com/CaraJ7/T2I-R1.
@@ 1 Introduction
  • Semantic-level CoT is the textual reasoning about the image to generate, which is introduced prior to the image generation. The semantic-level CoT designs the global structure of the image, e.g., the appearance and location of each object. In case the prompt requires reasoning shown in Fig. 2, the semantic-level CoT also helps to deduce the objects to generate. Optimizing the semantic-level CoT could explicitly manage the planning and reasoning of the prompt before the subsequent image tokens generation, making the generation easier.
@@
Figure 2: Visualization of the Image Generation Process of T2I-R1. All the prompts need reasoning or contain an uncommon scenario. We observe that T2I-R1 successfully deduces the true intention behind the prompt or provides a sensible imagination for the uncommon scenario (highlighted in the text) to produce a satisfying result compared with the baseline model, Janus-Pro.
@@
    Our resulting model, T2I-R1, that incorporates both levels of CoT using BiCoT-GRPO, demonstrates significant performance improvements and surpasses FLUX.1 across multiple established benchmarks.
@@ 3.1 Preliminary
$\mathcal{J}_{\text{GRPO}}(\theta)$ $=\mathbb{E}_{(q,a)\sim\mathcal{D},\{o_{i}\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}(\cdot\mid q)}$ (2)
@@ 3.3 BiCoT-GRPO
Then, we update the ULM by maximizing Equation 2. In practice, we incorporate the token-level policy gradient loss in [74], where the loss term is normalized over all the generated tokens to balance the reward on overly long semantic-level CoT.
@@ 3.4 Ensemble of Generation Rewards
Another option of the reward model is an object detector, e.g., GroundingDINO [38] and YOLO-world [8]. These open-vocabulary detection models accept an image along with object queries as input and output both the spatial positions and confidence scores for detected objects. This kind of vision expert serves as an ideal tool to evaluate the object’s existence and relationship concerning space and numbers. For implementation, we extract all objects ${obj_{i}}_{i=1}^{K}$ from the training image prompts, where $K$ represents the total number of objects. We then query the object detector to identify these objects within the generated image. For each object, we assign a binary existence score (1 if detected, 0 otherwise) and average these scores across all objects in the prompt. If the prompt contains a spatial relationship, we further leverage the detected location to validate its correctness. We calculate the relative distance and intersection over union (IoU) between the objects for the spatial score $\mathcal{R}_{\text{spatial}}$. If the number of the object $n_{obj_{i}}$ is specifically pointed out in the prompt, we compare the number with the detected number of the object $\hat{n}_{obj_{i}}$. The reward from the object detector $\mathcal{R}_{\text{Det}}$ is determined as:
@@
where $\mathcal{R}_{\text{spatial}}$ is 1 if the relative distance between the objects are larger than a threshold and the direction is right. If the direction is wrong, the reward is 0. Otherwise, we use the IoU as the spatial reward. We set $\alpha$ as 0.6 to encourage the correctness of the spatial relationship.
@@
The visual question answering (VQA) models are trained to answer questions based on the image input. The VQA models include earlier models prior to LLM, e.g., BLIP [33] and GIT [64], and LMMs like LLaVA [35]. We leverage these models to judge the existence and attributes of the objects. For example, if the image prompt is a red dog and a yellow cat, we first reformat each individual object $obj_{i}$ with its attribute as a question to the VQA model, i.e., a red dog? and a yellow cat?. Then, we record the probability for the model to answer Yes as $P_{\text{Yes}}^{i}$ and No as $P_{\text{No}}^{i}$. The reward for a prompt is calculated as:
$\mathcal{R}_{\text{VQA}}=\frac{1}{K}\sum_{i}\frac{P^{i}_{\text{Yes}}}{P_{\text{Yes}}^{i}+P_{\text{No}}^{i}}.$ (5)
@@
Lastly, we also employ the output reward model (ORM) proposed in [19] as a reward model. The ORM is fine-tuned from an LMM (e.g., LLaVA-OneVision [30]) specifically for evaluating the alignment between the prompt and the image. The fine-tuning is to instruct the model to output Yes if the image perfectly aligns with the image and output No otherwise. Therefore, we calculate $\mathcal{R}_{\text{ORM}}$ using the methodology similar to $\mathcal{R}_{\text{VQA}}$, except that we input the whole image prompt to the ORM instead of reformatting the prompt.
@@ 4 Experiment
4.1 Experimental Setup
@@
We test on T2I-CompBench [21] and WISE [44] to validate the effectiveness of our method. T2I-CompBench comprises 6,000 compositional text prompts evaluating three categories (attribute binding, object relationships, and complex compositions) and six sub-categories (color binding, shape binding, texture binding, spatial relationships, non-spatial relationships, and complex compositions). WISE consists of 1,000 text prompts spanning three categories (cultural common sense, spatial-temporal reasoning, and natural science) for evaluating world knowledge of the text-to-image models. To correctly generate an image, the model needs to reason about what the exact object or scenario is depicted in the prompt. We slightly modify the reasoning instruction on the WISE benchmark for more aligned results. We follow the official evaluation setting of the two benchmarks.
@@
[…]
4.2 Main Results
We compare T2I-R1 with leading text-to-image diffusion and autoregressive models on the T2I-CompBench and WISE benchmarks (in Table 1 and 2). We also provide the qualitative results in Fig. 5. Our method demonstrates substantial improvements over the baseline model, with average enhancements of 13% and 19% on T2I-CompBench and WISE, respectively. On T2I-CompBench, the most significant gains appear in attribute binding, with an average improvement of 19%. For the WISE benchmark, improvements are more evenly distributed across categories. When compared to the more powerful state-of-the-art diffusion models, T2I-R1 achieves superior or comparable results across both benchmarks. Notably, on T2I-CompBench, our method leads in five of six subtasks, with an exceptional performance in the spatial subtask (0.3378), surpassing previous SOTA results by over 5%. Similarly, for WISE, T2I-R1 excels in four of seven subtasks and achieves the highest overall score of 0.54, outperforming the robust FLUX.1-dev by 4%. Remarkably, our approach consistently achieves the leading results across all subtasks in both benchmarks when compared to other autoregressive models.
Remarkably, the improvement on T2I-Compbench benefits from the planning ability brought by the semantic-level CoT, which designs the complex scenarios before generation. While the enhancement of WISE is due to the reasoning capability from the semantic-level CoT, which deduces the true object or place depicted behind the prompt. The token-level CoT plays an important role in faithfully following the design to generate the image and ensure its visual appeal. As shown in Fig. 5, without the semantic-level CoT, the model fails to fully understand what the object or place is to generate, providing unaligned results. When lacking token-level CoT, we observe multiple artifacts in the images, showcasing relatively low image quality.
[…]
4.3 Reward Analysis
In this section, we experiment with the choice of reward functions and their combinations. We hope to provide some insights into how to choose the reward functions and combine them. Our results are shown in Table 3. We first experiment with the individual reward model. As shown in the table, HPM ($\mathtt{H}$) demonstrates superior performance in attribute binding but shows limited effectiveness in supervising object relationships, likely due to its weak relation comprehension capabilities. The object detector ($\mathtt{O}$) yields the least improvement in attribute binding among all other tested reward models, which aligns with expectations since our detector-based reward functions do not explicitly evaluate attributes. Any improvements observed stem solely from enhanced object existence ratios in the prompts. We observe that VQA model ($\mathtt{V}$) and ORM ($\mathtt{O}$) are both effective reward models with distinct strengths: the VQA model excels at improving attribute binding, while ORM demonstrates superior performance in object relationships. Then we experiment with multiple reward models. We start from the composition of HPM and object detector ($\mathtt{H+O}$), and progressively incorporate other reward models. Our findings indicate that both the HPM-object detector combination ($\mathtt{H+O}$) and the three-model integration of HPM, object detector, and VQA ($\mathtt{H+O+V}$) deliver balanced and satisfactory results across both attribute binding and relationship modeling tasks.
To obtain the optimal choice of reward models, we conduct a human study to evaluate the visual quality. Specifically, we select four options of reward models ($\mathtt{V}$, $\mathtt{O}$, $\mathtt{H+O}$, and $\mathtt{H+O+V}$) to generate an image from the same prompt. Then we ask humans to rank the four images and score them according to the rank (rank 1 for 3 points, rank 2 for 2 points, and so on). We randomly choose 30 prompts from each of the subtasks from the T2I-CompBench. The result is shown in the visual quality column in Table 3. We observe that ensemble rewards achieve better visual quality, with $\mathtt{H+O+V}$ obtaining slightly superior results. This improvement could be attributed to the implicit regularization provided by multiple rewards, preventing overfitting to a single reward model. Conversely, individual reward models fail to provide satisfactory quality despite high benchmark scores. To ensure visual appeal, we adopt the ensemble of three reward models ($\mathtt{H+O+V}$) for our final model.
[…]
4.4 Ablation Study
We validate the effectiveness of incorporating both semantic-level and token-level CoT. We first validate the effectiveness of the semantic-level CoT by comparing it with a baseline method that generates images using only the token-level CoT optimized with the GRPO method. This is the default text-to-image generation setting in Janus, whose result is shown in the third row in Table 4. Comparing the third and fourth row in the table, we find that semantic-level CoT generally brings performance improvements across both benchmarks tested. We witness a particularly significant gain on the WISE benchmark. This enhanced performance can be attributed to the textual reasoning capabilities inherent in semantic-level CoT. As illustrated in Fig. 5, our method could first clearly reason about the objects or phenomena described in the prompt through semantic-level CoT. This effectively decouples the reasoning and generation processes and thereby facilitates superior results. We also observe that training solely with token-level CoT substantially reduces the diversity of generated images, as demonstrated in Fig. 6. To quantify this effect, we evaluate image diversity by reusing the generated images from T2I-CompBench, where each prompt generates ten images. We compute the Vendi Score [15] across the ten images for each prompt. Results indicate that GRPO training without semantic-level CoT decreases the diversity score, whereas incorporating semantic-level CoT significantly improves diversity through varied textual planning.
@@ 4.4 Ablation Study
We also consider another situation: the semantic-level CoT is incorporated in image generation, but GRPO only optimizes the semantic-level CoT without the token-level CoT. This can be viewed as only enhancing the model’s high-level planning capabilities. The second row of Table 4 presents the result. The results show that optimizing semantic-level CoT exclusively yields smaller improvements compared to the joint optimization approach. Additionally, we find that optimizing both CoT types produces images with much better aesthetic quality compared with optimizing semantic-level CoT only. This indicates the necessity to jointly optimize both levels of CoT.
5 Conclusion

### v2

@@ Abstract
Recent advancements in large language models have demonstrated how chain-of-thought (CoT) and reinforcement learning (RL) can improve performance. However, applying such reasoning strategies to the visual generation domain remains largely unexplored. In this paper, we present T2I-R1, a novel reasoning-enhanced text-to-image generation model, powered by RL with a bi-level CoT reasoning process. Specifically, we identify two levels of CoT that can be utilized to enhance different stages of generation: (1) the semantic-level CoT for high-level planning of the prompt and (2) the token-level CoT for low-level pixel processing during patch-by-patch generation. To better coordinate these two levels of CoT, we introduce BiCoT-GRPO with an ensemble of generation rewards, which seamlessly optimizes both generated CoTs within the same training step. By applying our reasoning strategies to the baseline model, Janus-Pro, we achieve superior performance with 13% improvement on T2I-CompBench and 19% improvement on the WISE benchmark, even surpassing the state-of-the-art model FLUX.1. Code is available at: https://github.com/CaraJ7/T2I-R1.
@@ 1 Introduction
  • Semantic-level CoT is the textual reasoning about the image to generate, which is introduced prior to the image generation. The semantic-level CoT designs the global structure of the image, e.g., the appearance and location of objects. In case the prompt requires reasoning shown in Fig. 2, the semantic-level CoT also helps to deduce the objects to generate. Optimizing the semantic-level CoT could explicitly decouple the planning and reasoning of the prompt from the subsequent image tokens generation, making the generation easier.
@@
Figure 2: Visualization of the Image Generation Process of T2I-R1. All the prompts need reasoning or contain an uncommon scenario. We observe that T2I-R1 successfully deduces the true intention behind the prompt or provides a sensible imagination (highlighted in the text) to produce a satisfying result compared with the baseline model, Janus-Pro.
@@
    Our resulting model, T2I-R1, incorporates both levels of CoT using BiCoT-GRPO and demonstrates significant quantitative and qualitative improvements, surpassing FLUX.1 across multiple established benchmarks.
@@ 3.1 Preliminary
$\mathcal{J}_{\text{GRPO}}(\theta)$ $=\mathbb{E}_{(q,a)\sim\mathcal{D},\{o_{i}\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}(\cdot\mid q)}$
@@ 3.3 BiCoT-GRPO
Then, we update the ULM by maximizing Equation 3.1. In practice, we incorporate the token-level policy gradient loss in [83], where the loss term is normalized over all the generated tokens to balance the reward on overly long semantic-level CoT.
@@ 3.4 Ensemble of Generation Rewards
Another option of the reward model is an object detector, e.g., GroundingDINO [41] and YOLO-world [8]. These open-vocabulary detection models accept an image along with object queries as input and output both the spatial positions and confidence scores for detected objects. This kind of vision expert serves as an ideal tool to evaluate the object’s existence and relationship concerning space and numbers. For implementation, we extract all objects $\{obj_{i}\}_{i=1}^{K}$ from the training image prompts, where $K$ represents the total number of objects. We then query the object detector to identify these objects within the generated image. For each object, we assign a binary existence score (1 if detected, 0 otherwise) and average these scores across all objects in the prompt. If the prompt contains a spatial relationship, we further leverage the detected location to validate its correctness. We calculate the relative distance and intersection over union (IoU) between the objects for the spatial score $\mathcal{R}_{\text{spatial}}$. If the number of the object $n_{obj_{i}}$ is specifically pointed out in the prompt, we compare the number with the detected number of the object $\hat{n}_{obj_{i}}$. The reward from the object detector $\mathcal{R}_{\text{Det}}$ is determined as:
@@
where $\mathcal{R}_{\text{spatial}}$ is 1 if the relative distance between the objects is larger than a threshold and the direction is right. If the direction is wrong, the reward is 0. Otherwise, we use the IoU as the spatial reward. We set $\alpha$ as 0.6 to encourage the correctness of the spatial relationship.
@@
The visual question answering (VQA) models are trained to answer questions based on the image input. The VQA models include earlier models prior to LLM, e.g., BLIP [33] and GIT [70], and LMMs like LLaVA [38]. We leverage these models to judge the existence and attributes of the objects. For example, if the image prompt is a red dog and a yellow cat, we first reformat each individual object with its attribute as a question to the VQA model, i.e., a red dog? and a yellow cat?. Then, we record the probability for the model to answer Yes as $P_{\text{Yes}}^{i}$ and No as $P_{\text{No}}^{i}$. The reward for a prompt is calculated as: $\mathcal{R}_{\text{VQA}}=\frac{1}{K}\sum_{i}\frac{P^{i}_{\text{Yes}}}{P_{\text{Yes}}^{i}+P_{\text{No}}^{i}}.$
@@
Lastly, we also employ the output reward model (ORM) proposed in [19] as a reward model. The ORM is fine-tuned from an LMM (e.g., LLaVA-OneVision [30]) specifically for evaluating the alignment between the prompt and the image. The fine-tuning is to instruct the model to output Yes if the image perfectly aligns with the image and No otherwise. We calculate $\mathcal{R}_{\text{ORM}}$ using the methodology similar to $\mathcal{R}_{\text{VQA}}$, except that we input the whole image prompt to the ORM instead of reformatting the prompt.
@@ 4 Experiment
In this section, we first provide the main results of T2I-R1 in T2I-CompBench [22] and WISE [49] in Section 4.1. Then we present the results of different reward function combinations in Section 4.2 and the ablation study of the effectiveness of two levels of CoT in Section 4.3. Please refer to the Appendix A for more benchmark results (GenAI-Bench [36] and TIIF-Bench [73], detailed experiment setup, and more visualizations.
[…]
4.1 Main Results
We compare T2I-R1 with leading text-to-image diffusion and autoregressive models on the T2I-CompBench and WISE benchmarks (in Table 1 and 2). We also provide the qualitative results in Fig. 5. Our method demonstrates substantial improvements over the baseline model, with average enhancements of 13% and 19% on T2I-CompBench and WISE, respectively. On T2I-CompBench, the most significant gains appear in attribute binding, with an average improvement of 19%. For the WISE benchmark, improvements are more evenly distributed across categories. When compared to the more powerful state-of-the-art diffusion models, T2I-R1 achieves superior or comparable results across both benchmarks. Notably, on T2I-CompBench, our method leads in five of six subtasks, with an exceptional performance in the spatial subtask (0.3378), surpassing previous SOTA results by over 5%. Similarly, for WISE, T2I-R1 excels in four of seven subtasks and achieves the highest overall score of 0.54, outperforming the robust FLUX.1-dev by 4%. Remarkably, our approach consistently achieves the leading results across all subtasks in both benchmarks when compared to other autoregressive models. Remarkably, the improvement on T2I-Compbench benefits from the planning ability brought by the semantic-level CoT, which designs the complex scenarios before generation. While the enhancement of WISE is due to the reasoning capability from the semantic-level CoT, which deduces the true object or place depicted behind the prompt.
[…]
4.2 Reward Analysis
In this section, we experiment with the choice of reward functions and their combinations. We hope to provide some insights into how to choose the reward functions and combine them. Our results are shown in Table 3. We first experiment with the individual reward model. HPM ($\mathtt{H}$) demonstrates superior performance in attribute binding but shows limited effectiveness in object relationships, likely due to its weak relation comprehension capabilities. The object detector ($\mathtt{D}$) yields the least improvement in attribute binding, which aligns with expectations since our detector-based reward functions do not explicitly evaluate attributes. The improvements observed stem solely from enhanced object existence ratios in the prompts. We observe that VQA model ($\mathtt{V}$) and ORM ($\mathtt{O}$) are both effective reward models with distinct strengths: the VQA model excels at improving attribute binding, while ORM demonstrates superior performance in relationships. Then we experiment with multiple reward models. We start from the composition of HPM and object detector ($\mathtt{H+D}$), and progressively incorporate other reward models. Our findings indicate that both the HPM-object detector combination ($\mathtt{H+D}$) and the three-model integration of HPM, object detector, and VQA ($\mathtt{H+D+V}$) deliver balanced and satisfactory results in both attribute and relationship tasks. To obtain the optimal choice of reward models, we conduct a human study to evaluate the visual quality, detailed in Appendix B.3. We adopt the combination of the highest visual quality, the ensemble of three reward models ($\mathtt{H+D+V}$) for our final model.
[…]
4.3 Ablation Study
We validate the effectiveness of incorporating both semantic-level and token-level CoT. We first show the advantage of semantic-level CoT by comparing T2I-R1 with a baseline method that generates images using only token-level CoT optimized with GRPO. We witness a consistent gain on the benchmarks. However, we also find that training solely with token-level CoT substantially reduces image diversity, as demonstrated in Fig. 6 and 7. To quantify this, we provide the Vendi Score [15] as the diversity metric. We find that the diversity largely increases with semantic-level CoT while decreases without it. We then consider another setting to only optimize the semantic-level CoT to show the effectiveness of token-level CoT. The second row of Table 4 shows that optimizing semantic-level CoT exclusively yields smaller improvements compared to joint optimization. Additionally, optimizing both CoT types produces images with better aesthetic quality compared to optimizing semantic-level CoT only, as shown in Fig. 5. More details are in Appendix B.2.
5 Conclusion
Appendix A More Experiment Details
A.1 Experiment Setup
@@
We test on T2I-CompBench [22], WISE [49], GenAI-Bench [36], and TIIF-Bench [73] to validate the effectiveness of our method. T2I-CompBench comprises 6,000 compositional text prompts evaluating three categories (attribute binding, object relationships, and complex compositions) and six sub-categories (color binding, shape binding, texture binding, spatial relationships, non-spatial relationships, and complex compositions). WISE consists of 1,000 text prompts spanning three categories (cultural common sense, spatial-temporal reasoning, and natural science) for evaluating world knowledge of the text-to-image models. To correctly generate an image, the model needs to reason about what the exact object or scenario is depicted in the prompt. We slightly modify the reasoning instruction on the WISE benchmark for more aligned results. GenAI-Bench is a benchmark containing 1,600 complex, real-world text prompts collected from professional designers, which covers a broad spectrum of compositional text-to-visual generation elements, from basic aspects like scenes, attributes, and relationships to more professional ones, including counting, comparison, differentiation, and logical reasoning. TIIF-Bench is a comprehensive benchmark for fine-grained text-to-image model evaluation, featuring 36 novel prompt combinations across six compositional dimensions and 100 real-world designer-level prompts with rich aesthetic judgment. We follow the official evaluation setting of all the benchmarks.
Figure 7: More Visualization Result of the Image Diversity of a Single Prompt. We showcase the result of only token-level CoT optimized and both semantic-level and token-level CoT optimized.
Appendix B More Experiment Results
B.1 More Results
We provide the experiment results on GenAI-Bench in Table 5 and TIIF-Bench in Table 6. As shown in Table 5, T2I-R1 largely improves the baseline model, and in the meantime, achieves the highest overall score on both the basic and advanced prompts. Again, T2I-R1 surpasses FLUX.1 in both types of prompts and showcases a remarkable margin in the advanced prompt, probably attributed to the high-level reasoning capability granted by semantic-level CoT. We provide more qualitative examples in Fig. 8.
B.2 More Illustration of Ablation Study
To validate the effectiveness of the semantic-level CoT, we compare T2I-R1 with a baseline method that generates images using only the token-level CoT optimized with the GRPO method. This is the default text-to-image generation setting in Janus, whose result is shown in the third row in Table 4. Comparing the third and fourth row in the table, we find that semantic-level CoT generally brings performance improvements across both benchmarks tested. We witness a particularly significant gain on the WISE benchmark. This enhanced performance can be attributed to the textual reasoning capabilities inherent in semantic-level CoT. As illustrated in Fig. 5, our method could first clearly reason about the objects or phenomena described in the prompt through semantic-level CoT. This effectively decouples the reasoning and generation processes and thereby facilitates superior results. We also observe that training solely with token-level CoT substantially reduces the diversity of generated images, as demonstrated in Fig. 6, 7, 13, and 14. To quantify this effect, we evaluate image diversity by reusing the generated images from T2I-CompBench, where each prompt generates ten images. We compute the Vendi Score [15] across the ten images for each prompt. Results indicate that GRPO training without semantic-level CoT decreases the diversity score, whereas incorporating semantic-level CoT significantly improves diversity through varied textual planning.
Table 5: GenAI-Bench Evaluation Results. The best score is in blue, with the second-best score in green.
@@ B.2 More Illustration of Ablation Study
Table 6: TIIF-Bench Testmini Subset Evaluation Results. The best score is in blue, with the second-best score in green.
Model    Overall    Basic Following    Advanced Following    Designer          Avg    Attribute    Relation    Reasoning    Avg    Attribute +Relation    Attribute +Reasoning    Relation +Reasoning    Style    Text    Real World    short    long    short    long    short    long    short    long    short    long    short    long    short    long    short    long    short    long    short    long    short    long    short    long Llamagen [64]    41.67    38.22    53.00    50.00    48.33    42.33    59.57    60.32    51.07    47.32    35.89    32.61    38.82    31.57    40.84    47.22    49.59    46.22    46.67    33.33    0.00    0.00    39.73    35.62 LightGen [76]    53.22    43.41    66.58    47.91    55.83    47.33    74.82    45.82    69.07    50.57    46.74    41.53    62.44    40.82    61.71    50.47    50.34    45.34    53.33    53.33    0.00    6.83    50.92    50.55 Show-o [79]    59.72    58.86    73.08    75.83    74.83    79.83    78.82    78.32    65.57    69.32    53.67    50.38    60.95    56.82    68.59    68.96    66.46    56.22    63.33    66.67    3.83    2.83    55.02    50.92 Infinity [20]    62.07    62.32    73.08    75.41    74.33    76.83    72.82    77.57    72.07    71.82    56.64    54.98    60.44    55.57    74.22    64.71    60.22    59.71    80.00    73.33    10.83    23.83    54.28    56.89 Janus-Pro [7]    66.50    65.02    79.33    78.25    79.33    82.33    78.32    73.32    80.32    79.07    59.71    58.82    66.07    56.20    70.46    70.84    67.22    59.97    60.00    70.00    28.83    33.83    65.84    60.25 T2I-R1 (Ours)    68.59    67.19    82.90    81.63    86.50    83.00    83.47    79.43    78.73    82.46    69.05    68.00    71.64    69.47    72.43    69.95    69.40    70.40    60.00    63.33    27.60    26.24    67.54    60.45
@@
We also consider another situation to validate the effectiveness of token-level CoT: the semantic-level CoT is incorporated in the image generation process, as T2I-R1, but GRPO only optimizes the semantic-level CoT without the token-level CoT. This can be viewed as only enhancing the model’s high-level planning capabilities. The second row of Table 4 presents the result. The results show that optimizing semantic-level CoT exclusively yields smaller improvements compared to the joint optimization approach. Additionally, we find that optimizing both CoT types produces images with much better aesthetic quality compared with optimizing semantic-level CoT only. This indicates the necessity to jointly optimize both levels of CoT.
Figure 8: More Visualization Results. We provide the image generation results of the same prompt from four models: base model, the model with only semantic-level CoT optimized, the model with only token-level CoT optimized, and the model with both levels of CoT optimized.
Finally, we discuss the zero-shot potential of the baseline model to perform both semantic-level and token-level reasoning. Specifically, we apply the same image generation process of T2I-R1 directly to the baseline model, where the baseline model is first instructed to output the semantic-level CoT and then the token-level CoT. We term this method of generation as ‘Janus-Pro w/ zero-shot semantic-level CoT’ in Figure 9-12. As shown in the figure, zero-shot semantic-level CoT brings very marginal improvement, while T2I-R1 demonstrates a satisfying result. The reasons are twofold: (1) Zero-shot semantic-level CoT misses critical objects in the original prompt. As shown in Figure 12, the zero-shot semantic-level CoT misses the bird in the original prompt. (2) Zero-shot semantic-level CoT does not fit the model’s generation ability or provide useful information for generation. Although the semantic-level CoT in Figure 9-11 includes all the objects and relationships, the baseline model still fails to generate a satisfying result. This highlights the necessity of our proposed BiCoT-GRPO training method to build the synergy between the two levels of CoT and make them work together.
B.3 More Details about Reward Analysis
We conduct a human study to evaluate the visual quality of the generated images. Specifically, we select four options of reward models ($\mathtt{V}$, $\mathtt{O}$, $\mathtt{H+D}$, and $\mathtt{H+D+V}$) to generate an image from the same prompt. Then we ask humans to rank the four images and score them according to the rank (rank 1 for 3 points, rank 2 for 2 points, and so on). The humans are instructed to rank the images only based on the visual appeal. We employ eight graduate students to conduct the study to eliminate individual bias. We randomly choose 30 prompts from each of the subtasks from the T2I-CompBench. The result is shown in the visual quality column in Table 3. We observe that ensemble rewards achieve better visual quality, with $\mathtt{H+D+V}$ obtaining slightly superior results. This improvement could be attributed to the implicit regularization provided by multiple rewards, preventing overfitting to a single reward model. Conversely, individual reward models fail to provide satisfactory quality despite high benchmark scores.
B.4 Hyperparameters
All of our experiments are conducted on 8 H800. Our training procedure lasts about 16 hours. We provide the detailed training hyperparameters in Table 7.
Table 7: T2I-R1 training hyperparameters.
Name
Learning rate 1e-6
Beta $\beta$ 0.01
Group Size $G$ 8
Classifier-Free Guidance Scale 5
Max Gradient Norm 1.0
Batchsize 8
Training Steps 1,600
Gradient Accumulation Steps 2
Image Resolution $h\times w$ $384\times 384$
Figure 9: Visualization Results of Semantic-level CoT. We provide the image generation results of the same prompt from three settings: base model, base model with zero-shot semantic-level CoT, and T2I-R1. For the setting of base model with zero-shot semantic-level CoT, we use the same generation pipeline of T2I-R1 directly on the base model. We employ the same prompt of T2I-R1 to instruct the base model to generate a zero-shot semantic-level CoT, which we visualize in the figure and provide a comparison of the semantic-level CoT generated by T2I-R1.
Figure 10: Visualization Results of Semantic-level CoT. We provide the image generation results of the same prompt from three settings: base model, base model with zero-shot semantic-level CoT, and T2I-R1. For the setting of base model with zero-shot semantic-level CoT, we use the same generation pipeline of T2I-R1 directly on the base model. We employ the same prompt of T2I-R1 to instruct the base model to generate a zero-shot semantic-level CoT, which we visualize in the figure and provide a comparison of the semantic-level CoT generated by T2I-R1.
Figure 11: Visualization Results of Semantic-level CoT. We provide the image generation results of the same prompt from three settings: base model, base model with zero-shot semantic-level CoT, and T2I-R1. For the setting of base model with zero-shot semantic-level CoT, we use the same generation pipeline of T2I-R1 directly on the base model. We employ the same prompt of T2I-R1 to instruct the base model to generate a zero-shot semantic-level CoT, which we visualize in the figure and provide a comparison of the semantic-level CoT generated by T2I-R1.
Figure 12: Visualization Results of Semantic-level CoT. We provide the image generation results of the same prompt from three settings: base model, base model with zero-shot semantic-level CoT, and T2I-R1. For the setting of base model with zero-shot semantic-level CoT, we use the same generation pipeline of T2I-R1 directly on the base model. We employ the same prompt of T2I-R1 to instruct the base model to generate a zero-shot semantic-level CoT, which we visualize in the figure and provide a comparison of the semantic-level CoT generated by T2I-R1.
Figure 13: More Visualization Result of the Image Diversity of a Single Prompt. We showcase the result of the baseline model, Janus-Pro, and T2I-R1.
Figure 14: More Visualization Result of the Image Diversity of a Single Prompt. We showcase the result of the baseline model, Janus-Pro, and T2I-R1.
Appendix C Limitations and Future Work
While this work explores the text-to-image generation task, it requires more exploration on how to apply this paradigm to video generation tasks. Video generation tasks are more complex regarding the reward design and the base model. For the reward design, how to apply dense rewards on each generated frame is still an open question. Besides, there exists no understanding and generation unified model for videos, so BiCoT-GRPO cannot be used directly. Meanwhile, the current inference time of video generation is too long for the current GRPO paradigm. How to balance the training time and effect needs further study.